OpenStudy (anonymous):

What is the slope-intercept form equation of the line that passes through (2, 4) and (4, 10)?

OpenStudy (alexandervonhumboldt2):

first we need to count slope using slope formula: (y_1-y_2)/(x_1-x_2)=(4-10)/(2-4)=(-6)/(-2)=3 slope is 3. now substitute any coordinate into the equation: 4=2*3+b b=-2 thus the equation is y=3x-2